# On the Existence of Steady-State Solutions to the Equations Governing Fluid Flow in Networks

### 23 Jan 2024 v3

@@ 3 Background
If $\mathrm{deg}(\bm{f},\mathcal{D},\bm{0})\neq 0$, then $\bm{f}$ must have a zero/root in $\mathcal{D}$.
Proposition 4 (Homotopy invariance).
In order to use Proposition 4, first we recognize that the solution of system  (4d) and (5d) may be thought of as a zero or root of a non-linear operator on $\mathbb{R}^{|E|+|V|}$. In order to conform to Proposition 4, we need to modify system (4d) or (5d) to introduce a scalar parameter $s\in[0,1]$ such that the operator has the form $F:\mathbb{R}^{|E|+|V|}\times\mathbb{R}\longrightarrow\mathbb{R}^{|E|+|V|}$. This is easily accomplished in system (7d) so that setting $s=1$ recovers system (5d) as follows:
@@
The following result (restated with slight modification from [1]) that guarantees the invertibility of a linear system with particular structure is essential for the subsequent analysis.
Proposition 5.
$a^{*}_{i}x_{i}-a^{*}_{j}x_{j}-b^{*}_{ij}y_{ij}=0\quad\forall(i,j)\in E,$ (8a)
$x_{i}=0\quad\forall i\;\in N_{s}.$ (8c)
4 Results and Implications
@@
A unique solution to $\mathcal{G}\mathcal{F}_{\pi}$ system (5d) exists.
@@
We shall appeal to Proposition 4 to assert the existence of a solution. Accordingly, we first observe that $F((\bm{\phi},\bm{\pi}),0)$ is a linear system of the same form as Proposition 5 and hence has a unique solution in the open hypercube $\mathcal{D}_{\mathcal{H}}$ in (6) constructed in Proposition 2. Moreover, $F((\bm{\phi},\bm{\pi}),s)$ is continuous on $\mathrm{cl}(\mathcal{D}_{\mathcal{H}})$ and by construction no solution, if it exists, can lie on the boundary of the hypercube. Hence by Proposition 4, $F((\bm{\phi},\bm{\pi}),s)$ has a solution in $\mathcal{D}_{\mathcal{H}}$ for every $s\in[0,1]$. Thus, for $s=1$, we conclude that the system (5d) has a generalized-potential solution. Once existence is assured, uniqueness follows from Lemma 2. ∎
@@
If the (unique) solution to (5d) has values of the potential $\pi_{i}\in\mathbb{R}\;\forall i\in V$ such that they can be inverted to find pressures $p_{i}\in\mathbb{R}$, such that $\pi(p_{i})=\pi_{i}$, then a generalized pressure solution exists for system (4d).
We note in passing that for $\pi^{\mathrm{ideal}}$ (2a), inversion is possible only for non-negative values but since the expression for $\pi^{\mathrm{cnga}}$ (2b) is a cubic polynomial, a preimage can be found corresponding to any real value of the potential.
4.1 Ideal Gas
Theorem 1 and Corollary 1 have far reaching implications, laying to rest all ambiguity about the solvability of the steady-state gas flow equations on networks for the case of the ideal gas. They assure us that steady-state network flow for an ideal gas (corresponding to system (5d)) always has a unique, generalized potential solution. Furthermore, the expression for $\pi$ in (2a) makes it clear that a generalized-pressure solution can be derived contingent on non-negative potentials.
It was pointed out in Remark 2 that our definition of generalized solution allowed the potentials, pressures and flows to take any real values. However, at this juncture, we can allow physical considerations to make a distinction between the equations being solvable and feasible.
As outlined in [1], for both the ideal and the non-ideal gas (with CNGA equation of state), a feasible solution will have non-negative pressures and non-negative flow in compressors. Otherwise, the solution/system is infeasible.
In light of this statement, one may qualify the questions (Q1), (Q2) asked earlier to now read:
@@ 4.1 Ideal Gas
  2. (Q5)
    For a non-ideal gas (with the CNGA equation of state), does the system (4d) have a feasible solution?
@@
4.2 Non-ideal Gas
@@
Figure 1: The plots above show the magnitude and distribution of the relative error $\mathrm{e}_{\mathrm{rel}}(\alpha,p)$ in (10) for two different choices of $g(\alpha)$. The colour bars demonstrate how the error is reduced drastically by choosing the optimal $g(\alpha)$ in (9). With this choice, we may write $\pi(\alpha p)\approx\gamma_{\mathrm{opt}}\pi(p)$ so that the system (4d) for a non-ideal gas is approximated by (5d).
We have thus far made no assertions about the existence of steady-state solutions when the gas obeys the CNGA equation of state and thus no attempt to answer (Q2) and (Q5) articulated earlier. For the existence of solutions to system (4d) for a non-ideal gas, it is tempting to construct a continuous homotopy function $\chi(p,s)$ that satisfies $\chi(p,0)=\pi^{\mathrm{ideal}}(p)$ and $\chi(p,1)=\pi^{\mathrm{cnga}}(p)$ to replace $\pi(p)$ in system (4d) and reuse the artifice of Proposition 4 to analyse the system (4d). However, it is not clear how to estimate an upper bound for the solution due to the term $\chi(p,s)$ analogous to Proposition 2; indeed we cannot be sure that no solution will lie on the prescribed boundary for any $s\in[0,1]$.
In abandoning this approach and conceding the futility of the task associated with (Q2) and (Q5), it is natural to consider if the theory developed for the ideal gas could prove useful in some other way for the non-ideal gas. Towards this end, we look to examine the possibility of approximating system (4d) by (5d) for the CNGA equation of state. We already know that because the potential function for the CNGA equation of state (2b) violates Lemma 1, (5b) does not follow from (4b).
However, if we could write $\pi(\alpha p)\approx g(\alpha)\pi(p)$ for some function $g$, then (5d) could be a system that approximates (4d) for the CNGA equation of state in the sense that the solution to (5d) is an acceptable approximant to the solution of (4d) when the latter exists. To this end, we could look to determine
@@
such that the residual $\pi(\alpha p)-g(\alpha)\pi(p)$ is minimized in a least-squares sense implied by the quantity
@@
for $g(\alpha)=\gamma_{\mathrm{opt}}$ with that of $g(\alpha)=\alpha^{2}$ (which would be applicable for the ideal gas (2a)) in Figure 1. It can be seen that the error is reduced drastically by choosing the optimal $g(\alpha)$ in (9). With this choice, we may write $\pi(\alpha p)\approx\gamma_{\mathrm{opt}}\pi(p)$ so that the system (4d) for a non-ideal gas is approximated by (5d).
Let $\bm{x}_{a}$ be the pressure-flux pair obtained from inversion of the potentials in the solution to system (5d) with the choice $\gamma_{\mathrm{opt}}$ in (5b). Note that $\bm{x}_{a}$ can always be found since a unique generalized potential solution to system (5d) always exists, and a pressure corresponding to a potential can always be found for the cubic polynomial (2b). Suppose that $\bm{x}$ is the solution to the original system (4d). If we denote by $\bm{R}(\bm{y})$ the vector of residuals of the system (4d) evaluated at $\bm{y}$, then $||\bm{R}(\bm{x})||_{\infty}\approx 0$.
Motivated by standard ideas in inversion of linear systems [17], we want an approximate solution to have a small residual, and also be close to the true solution. Thus, the two quantities that measure the acceptability of $\bm{x}_{a}$ as an approximation of $\bm{x}$ are
$\dfrac{||\bm{x}_{a}-\bm{x}||_{\infty}}{||\bm{x}||_{\infty}}\text{ and }||\bm{R}(\bm{x}_{a})||_{\infty}.$
Unfortunately, we are unable to estimate these quantities theoretically. The only qualitative assertion we can make is that the solution $\bm{x}_{a}$ ought to satisfy (4a), (4d) and that the quantity $||\bm{R}(\bm{x}_{a})||_{\infty}$ will be determined by (4b) alone. We also expect that a good approximate solution will serve as a good initial guess in a Newton-Raphson algorithm and lead to convergence within a handful of iterations.
Figure 2: Schematic of the GasLib-40 network. The junctions that connect compressors are shown in red and the slack junction is shown in blue.
@@
While no claim can be made regarding the existence of solutions to system (4d) in general, these numerical experiments support our assertion that the solution $\bm{x}_{a}$ of system (5d) is a good approximation to the computed solution $\bm{x}$ of system (4d). Moreover, starting with the initial guess $\bm{x}_{a}$, the convergence to $\bm{x}$ was attained in exactly 2 Newton iterations in each instance.
The comparison between the solutions for an ideal gas and a non-ideal gas has already been analyzed in [1] and is hence not considered here.
Figure 3: Cumulative distribution function for GasLib-40 showing the residual $O(10^{-2})$ as well as the relative error $O(10^{-3})$ of the approximate solution for the non-ideal gas.
4.3 Other fluids
The results obtained so far can be generalized to other fluid flow networks once the equation of state that describes the pressure-density relationship is specified. For instance, in water distribution networks or in carbon capture and sequestration, the incompressibility of water and carbon-dioxide implies that the equation of state is $\rho(p)=\rho^{*}$ (a constant), which makes the potential function linear (and simpler) [19] in (1). The other difference is the use of pumping devices which provide a given additive increase in potential or pressure and simplify the analysis [20] in comparison to a given multiplicative increase as in (4b) or (5b) considered for compressors.
@@ 5 Conclusion
The arguments and proofs presented in this letter lay to rest the question of existence of solutions to potential-driven, steady-state fluid flow in a network. As the question of existence of solutions had in it the embedded complexity of the network topology and the equation of state of the gas, the results are also applicable to other networks carrying fluids such as water [19] or carbon-dioxide. Importantly, it gives support to computational work by enabling correct diagnosis of algorithmic failure, problem stiffness and non-convergence by answering the question of existence of solutions. Future work could look towards determining solutions by homotopic continuation methods [12] starting with a less complicated system with a guaranteed solution as exemplified by [15].

### 08 Apr 2024 v4

@@ 4 Results and Implications
$a^{*}_{i}x_{i}-a^{*}_{j}x_{j}-b^{*}_{ij}y_{ij}=0\quad\forall(i,j)\in E,$ (8a)
$x_{i}=0\quad\forall i\;\in N_{s}.$ (8c)
In order to invoke Proposition 2, we interpret the solution of system  (6d) as a zero or root of a nonlinear operator on $\mathbb{R}^{|E|+|V|}$ and modify system (6d) to introduce a scalar parameter $s\in[0,1]$ such that the operator now (9d) has the form
$F:\mathbb{R}^{|E|+|V|}\times\mathbb{R}\longrightarrow\mathbb{R}^{|E|+|V|}.$
Setting $s=1$ in system (9d) recovers system (6d).
@@
A solution exists for $\mathcal{G}\mathcal{F}_{\pi}$ system (6d) and it is unique.
@@
Observe that $F((\bm{\phi},\bm{\pi}),0)$ is a linear system of the same form as Proposition 5 and hence has a unique solution in the open hypercube $\mathcal{D}_{\mathcal{H}}$ in (7) constructed in Proposition 4. Moreover, $F((\bm{\phi},\bm{\pi}),s)$ is continuous on $\mathrm{cl}(\mathcal{D}_{\mathcal{H}})$ and by construction no solution can lie on the boundary of the hypercube. Hence by Proposition 2, $F((\bm{\phi},\bm{\pi}),s)$ has a solution in $\mathcal{D}_{\mathcal{H}}$ for every $s\in[0,1]$. For $s=1$, the system (6d) then has a generalized-potential solution. Uniqueness follows from Lemma 2. ∎
@@
If the (unique) generalized-potential solution to (6d) has values of the potential $\pi_{i}\in\mathbb{R}\;\forall i\in V$ such that they can be inverted to find pressures $p_{i}\in\mathbb{R}$ satisfying $\pi(p_{i})=\pi_{i}$, then a generalized-pressure solution exists for system (5d).
We note in passing that while inversion is possible only for non-negative values of $\pi^{\mathrm{ideal}}$ (4a), for $\pi^{\mathrm{cnga}}$ (4b) a preimage can be found corresponding to any real value of the potential as the expression is a cubic polynomial.
4.1 Ideal Gas Equation of State
Theorem 1 and Corollary 1 have far-reaching implications, laying to rest all ambiguity about the solvability of the steady-state gas flow equations on networks for the case of the ideal gas. The steady-state system for an ideal gas (corresponding to system (6d)) always has a unique, generalized potential solution. Furthermore, a generalized-pressure solution can be derived contingent on non-negative potentials. In light of this statement, one may qualify the question (Q1) asked earlier to now read:
@@ 4.1 Ideal Gas Equation of State
4.2 Non-ideal Gas with CNGA Equation of State
@@
Figure 1: The plots above show the magnitude and distribution of the relative error $\mathrm{e}_{\mathrm{rel}}(\alpha,p)$ in (11) for two different choices of $g(\alpha)$ for the CNGA equation of state. The colour bars demonstrate how the error is reduced drastically by choosing the optimal $g(\alpha)$ in (10). With this choice, we may write $\pi(\alpha p)\approx\gamma_{\mathrm{opt}}\pi(p)$ so that the system (5d) for the CNGA equation of state is approximated by (6d).
We have thus far made no attempt to answer (Q2) articulated earlier because we already know generalized-pressure solutions may not exist [1]. However, it is still of interest in the context of $\mathcal{G}\mathcal{F}_{p}$ system (5d), to construct a continuous homotopy function $\chi(p,s)$ that satisfies $\chi(p,0)=\pi^{\mathrm{ideal}}(p)$ and $\chi(p,1)=\pi^{\mathrm{cnga}}(p)$ to replace $\pi(p)$ in system (5d) and examine why Proposition 2 fails to apply. The failure is owing to the term $\chi(p,s)$, where it is not possible to estimate an upper bound for the solution analogous to Proposition 4. Thus we are no closer to any kind of answer regarding existence of solution for the CNGA equation of state in (Q2). However, it is natural to consider if the theory developed could prove useful in some other way.
Even though (5b) does not imply (6b) for $\pi^{\mathrm{cnga}}$ (4b), can we approximate $\pi(\alpha p)\approx g(\alpha)\pi(p)$ for some function $g$, so that $\mathcal{G}\mathcal{F}_{\pi}$ (6d) approximates $\mathcal{G}\mathcal{F}_{p}$ (5d) for the CNGA equation of state ? This approximation will be sensible only if the solution to $\mathcal{G}\mathcal{F}_{\pi}$ (6d) (which always exists by Theorem 1) is an acceptable approximant to the solution of $\mathcal{G}\mathcal{F}_{p}$ (5d) (when the latter exists). To investigate if that is so, we look to determine
@@
for the CNGA equation of state such that the residual $\pi(\alpha p)-g(\alpha)\pi(p)$ is minimized in a least-squares sense implied by the quantity
@@
for $g(\alpha)=\gamma_{\mathrm{opt}}$ with that of the choice $g(\alpha)=\alpha^{2}$ in Figure 1. It can be seen that the error is reduced drastically by choosing the optimal $g(\alpha)$ in (10). With this choice, we may write $\pi(\alpha p)\approx\gamma_{\mathrm{opt}}\pi(p)$ so that the system $\mathcal{G}\mathcal{F}_{p}$ (5d) for the CNGA equation of state is approximated by $\mathcal{G}\mathcal{F}_{\pi}$ (6d).
Let $\bm{x}_{a}$ be the pressure-flux pair obtained from inversion of the potentials in the solution to $\mathcal{G}\mathcal{F}_{\pi}$ system (6d) with the choice $\gamma_{\mathrm{opt}}$ in (6b). Note that $\bm{x}_{a}$ can always be found since a unique generalized potential solution to $\mathcal{G}\mathcal{F}_{\pi}$ system (6d) always exists, and a pressure corresponding to a potential can always be found for the cubic polynomial (4b). We now compute a solution $\bm{x}$ (assuming it exists) to the original $\mathcal{G}\mathcal{F}_{p}$ system (5d). If we denote by $\bm{R}(\bm{y})$ the vector of residuals of the $\mathcal{G}\mathcal{F}_{p}$ system (5d) evaluated at $\bm{y}$, then $||\bm{R}(\bm{x})||_{\infty}\approx 0$.
Motivated by standard ideas in the inversion of linear systems [16], an approximate solution should have a small residual and also be close to the true solution. Thus, the two quantities that measure the acceptability of $\bm{x}_{a}$ as an approximation of $\bm{x}$ are the relative error and the residual expressed by
$\dfrac{||\bm{x}_{a}-\bm{x}||_{\infty}}{||\bm{x}||_{\infty}}\text{and }||\bm{R}(\bm{x}_{a})||_{\infty}.$ (12)
The only qualitative assertion we can make is that the solution $\bm{x}_{a}$ ought to satisfy (5a), (5d) and that the quantity $||\bm{R}(\bm{x}_{a})||_{\infty}$ will be determined by (5b) alone. We also expect that a good approximate solution will serve as a good initial guess in a Newton-Raphson algorithm and lead to convergence within a handful of iterations.
@@
These numerical experiments support our assertion that the solution $\bm{x}_{a}$ of system (6d) is a good approximation to the computed solution $\bm{x}$ of system (5d) whenever the latter exists. Moreover, starting with the initial guess $\bm{x}_{a}$, the convergence to $\bm{x}$ was attained in exactly 2 Newton iterations in each instance. The comparison between the solutions for an ideal gas and a non-ideal gas has already been analyzed in [1] and is hence not considered here.
Figure 2: Cumulative distribution function for GasLib-40 showing the residual $O(10^{-2})$ (red) as well as the relative error $O(10^{-3})$ (blue) defined by (12) for the non-ideal gas.
4.3 Generalizations
The results obtained so far can be generalized to flow networks with different fluids. The three points of difference can come from the equation of state describing the fluid, the model of the pumping device, and the viscous resistance function for the fluid (in this letter, the resistance function was $\phi|\phi|$).
In general, for any equation of state that has the form of a scaled monomial, the existence result in Theorem 1 remains unchanged. This is so in water distribution networks or in carbon capture and sequestration where the incompressibility of water and carbon dioxide implies that the equation of state is $\rho(p)=\rho^{*}$ (a constant) [18]. The difference in the model of pumping devices could manifest in a given additive increase in potential or pressure [19] in place of a given multiplicative increase considered in (5b) or (6b), but Theorem 1 still holds. Finally, Theorem 1 continues to hold as long as the viscous resistance function chosen is monotonic and differentiable.
4.4 Use of Existence Result in Optimization
We shall now illustrate how the existence of a unique solution to $\mathcal{G}\mathcal{F}_{\pi}$ can be used to detect infeasibilities in the so-called optimal gas flow problem (OGF) that is solved regularly by natural gas pipeline network operators.
The steady-state OGF [3] is a nonlinear optimization problem described as follows: Given (vectors of) non-slack base nodal injections $\tilde{\bm{q}}$, slack nodal pressures $\tilde{\bm{p}}$, the minimum pressure value $p^{\min}$ at any non-slack node, and bounds ($\alpha^{\min}$ and $\alpha^{\max}$) for the compressor ratios, determine the vector of compressor ratios $\bm{\alpha}$ to minimize the total compressor power [15] while satisfying the steady-state gas flow physics $\mathcal{G}\mathcal{F}_{p}$ Eq. (5d) with $\bm{q}^{*}=\tilde{w}\cdot\tilde{\bm{q}}$, $\bm{p}^{*}=\tilde{\bm{p}}$, and $\bm{\alpha}^{*}=\bm{\alpha}$. Here a given injection scaling factor $\tilde{w}\geqslant 1$ is used to parametrize the OGF.
Mathematically, the OGF is formulated as:
$\mathrm{OGF}(\tilde{w}):$ $\min\sum_{(i,j)\in C}\phi_{ij}(\alpha_{ij}^{m}-1)\text{ subject to: }$: (13a)
$\mathrm{OGF}(\tilde{w}):$ $~{}~{}\mathcal{G}\mathcal{F}_{p}\text{ with ($\bm{\alpha}$, $\tilde{\bm{p}}$, $\tilde{w}\cdot\tilde{\bm{q}}$), }$, (13b)
$\mathrm{OGF}(\tilde{w}):$ $~{}~{}\alpha^{\min}\leqslant\alpha_{ij}\leqslant\alpha^{\max}~{}\forall(i,j)\in C,$ (13c)
$\mathrm{OGF}(\tilde{w}):$ $~{}~{}\tilde{p}_{i}\geqslant p^{\min}~{}\forall i\in N_{ns}.$ (13d)
Eventually, $\tilde{\bm{q}}$ scaled by a large enough $\tilde{w}$ will cause withdrawals at some nodes large enough to lower nodal pressures below $p^{\min}$ even if compression ratios are at their maximum. Thus, if we adopt the following procedure – (i) Set $\alpha_{ij}$ = $\alpha^{\max}$, (ii) Solve $\mathcal{G}\mathcal{F}_{\pi}$ (By Theorem 1, unique solution exists), we can guarantee that the $\mathrm{OGF}(\tilde{w})$ is infeasible if for any $i\in N_{ns}$, either $\pi_{i}<0$ or $p_{i}<p^{\min}$. For GasLib-40, this procedure shows that $\mathrm{OGF}(\tilde{w})$ is infeasible for any $\tilde{w}\geqslant 2.135$ given $\alpha^{\min}=1$, $\alpha^{\max}=1.5$, and $p^{\min}=300~{}$\mathrm{p}\mathrm{s}\mathrm{i}$$.
There are no general techniques to provide infeasibility guarantees for nonlinear optimization problems and the infeasibility guarantees provided in this section was made possible solely due to the existence result for $\mathcal{G}\mathcal{F}_{\pi}$ derived in Theorem 1.
@@ 5 Conclusion
The arguments and proofs presented in this letter lay to rest the question of existence of solutions to potential-driven, steady-state fluid flow in a network when the potential functions are scaled monomials (Lemma 1) such as for the ideal gas. However, no conclusion can be drawn regarding existence of solutions for potential functions that do not obey Lemma 1. However, for gases following the CNGA equation of state, while the question of existence remains open, we construct an alternative system that always has a unique solution and show that the solution to this system is a good approximant of the true solution. The results are also applicable to other networks carrying fluids such as water [18] or carbon dioxide. Importantly, it gives support to computational work by enabling correct diagnosis of algorithmic failure, problem stiffness and non-convergence. Future work could look towards determining solutions by homotopic continuation methods [11] starting with a less complicated system with a guaranteed solution as exemplified by [13].